Jeff and Diana are working on Math homework. Jeff can comlete 4 math problems in 11 minutes. Diana can complete 7 math problems in 25 minutes. Who is faster?
11/4=25/7
2.75=3.57
Jeff is faster.
Is this right

Found 3 solutions by Theo, Alan3354, palanisamy:
Answer by Theo(13342)

jeff completes 4 in 11 minutes
dana completes 7 in 25 minutes.

jeff takes 11/4 = 2.75 minutes to complete 1 problem.

dana takes 25/7 = 3.57 minutes to complete 1 problem.

jeff is faster.

you are correct.

another way of looking at is rate * time = units produced.

with jeff, you get r * 11 = 4

with dana, you get 4 * 25 = 7

jeff's rate is 4/11 = .36 problems per minute.

dana's rate is 7/25 = .28 problems per minute.

jeff's rate per minute is higher than dana's.

jeff is faster.


Answer by Alan3354(69443)

Jeff and Diana are working on Math homework. Jeff can comlete 4 math problems in 11 minutes. Diana can complete 7 math problems in 25 minutes. Who is faster?
---------------
Faster is the most problems per minute.
4/11 vs 7/25
4/11 = 0.36 problems/min
7/25 = 0.28 problems/min
-------
Jeff does more problems per minute.
--------------------
You can find the common denominator, too
4/11 = 100/275
7/25 = 77/275
100 is more than 77
----------------
The inverse is minutes per problem, then the smaller number is faster.
11/4 = 2.75 mins per problem
25/7 =~ 3.57 mins per problem
--------------------
Maybe Jeff did the easy problems.

Answer by palanisamy(496)

Jeff can complete 4 math problems in 11 minutes
In one minute Jeff can complete 4/11 = 0.3636 part of a problem.
Next, Diana can complete 7 math problems in 25 minutes.
In one minute Diana can complete 7/25 = 0.28 part of a problem.
So Jeff is faster